# Automating Knowledge Discovery from Scientific Literature via LLMs: A Dual-Agent Approach with Progressive Ontology Prompting

### v1

@@ Abstract
To address the challenge of automating knowledge discovery from a vast volume of literature, in this paper, we introduce a novel framework based on large language models (LLMs) that combines a progressive ontology prompting (POP) algorithm with a dual-agent system, named LLM-Duo, designed to enhance the automation of knowledge extraction from scientific articles. The POP algorithm utilizes a prioritized breadth-first search (BFS) across a predefined ontology to generate structured prompt templates and action orders, thereby guiding LLMs to discover knowledge in an automatic manner. Additionally, our LLM-Duo employs two specialized LLM agents: an explorer and an evaluator. These two agents work collaboratively and adversarially to enhance the reliability of the discovery and annotation processes. Experiments demonstrate that our method outperforms advanced baselines, enabling more accurate and complete annotations. To validate the effectiveness of our method in real-world scenarios, we employ our method in a case study of speech-language intervention discovery. Our method identifies 2,421 interventions from 64,177 research articles in the speech-language therapy domain. We curate these findings into a publicly accessible intervention knowledge base that holds significant potential to benefit the speech-language therapy community.
footnotetext: Preprint. Under Review.
Introduction
@@
With millions of research articles published annually, the overwhelming amount of existing scientific knowledge poses extreme challenges and opportunities for researchers to access knowledge through advanced analytical tools and interdisciplinary approaches. Knowledge discovery from scientific literature enables researchers to keep abreast of the latest developments in their domain and facilitate valuable insights that can significantly enhance their work (Usai et al. 2018; Wang et al. 2023). However, in such a vast ocean of data, only a very limited fraction of knowledge is collected and curated due to the low efficiency of the human review process. For example, in healthcare, evidence-based intervention refers to practices and treatments that are grounded in systematic research and have been proven effective through controlled studies (Rutten et al. 2021; Melnyk and Fineout-Overholt 2022). It emphasizes the use of evidence from well-designed and well-conducted research as the foundation for healthcare decision-making (Sackett 1997). One of the biggest challenges for healthcare providers is efficiently identifying relevant intervention evidence from research articles, thus underscoring the necessity of automating knowledge discovery.
Advancements in large language models (LLMs) present a significant opportunity for automating knowledge discovery in scientific literature. LLMs have been employed to categorize research papers, extract key findings, summarize complex studies, and streamline the review process effectively (Achiam et al. 2023; Guo et al. 2023). Some recent research (Li et al. 2023; Kim et al. 2024) utilize LLMs in conjunction with human annotators to reduce the human annotation burden. However, facing such a vast volume of domain knowledge, automating knowledge discovery is necessary to process and analyze information efficiently. For automating knowledge discovery from literature with LLMs, one big challenge is the limited context window length of LLMs (Wadhwa, Amir, and Wallace 2023). This constraint restricts the amount of input text that the models can process at one time, potentially leading to incomplete analysis and missed connections between data points spread across larger documents. To address this issue, the Retrievel-Augmented Generation (RAG) technique (Lewis et al. 2020; Gao et al. 2023) can be instrumental by combining a powerful retrieval component with a generative model, allowing the system to access a broader corpus of information beyond the immediate context window of a single model.
@@
In this paper, we address the challenge of automating knowledge discovery through LLMs and formulate it as a problem of prompt design and scheduling based on a predefined ontology graph structure. To tackle this problem, we introduce a progressive ontology prompting (POP) algorithm that employs an outdegree-prioritized breadth-first search (BFS) across the ontology graph to create a series of prompt templates and action orders to guide LLMs. Additionally, to enhance the annotation quality, we further propose LLM-Duo, an interactive annotation framework for domain-specific knowledge discovery by leveraging the power of LLMs while addressing the limitations of LLMs. Specifically, this framework integrates two LLM agents working both collaboratively and adversarially to improve the discovery quality: 1) explorer, a chatbot based on the RAG technique, working in a zero-shot setting to generate annotation results while arguing with the evaluator to justify its answers; and 2) evaluator, an LLM tasked with evaluating annotation and providing feedback to assist explorer to refine its annotations.
@@
To demonstrate the effectiveness of our approach in practice, we apply our method in a practical scenario, using speech-language intervention discovery as a case study. In our experiments, we compare our annotation framework to a number of advanced baselines including long context LLM (i.e., GPT-4-Turbo with 128k context window length), and RAG-based chatbot with prompting methods including Chain-of-Thought (CoT) (Wei et al. 2022) and SelfRefine (Madaan et al. 2024). The results show that our framework achieves more accurate and complete discoveries over baselines. The main contributions of our work are outlined as follows:
  • We formulate automated knowledge discovery with LLMs as a prompt design and scheduling problem, design a novel POP algorithm to transform a knowlege graph (KG) ontology into structured prompts and action orders, and achieve automatic knowledge discovery from literature.
  • We propose a novel annotation framework with two LLM agents that collaborate and compete to refine knowledge findings, which achieves superior performance over advanced baselines.
  • We apply our method to speech-language intervention discovery and successfully curate an intervention knowledge base that can serve as a valuable resource for the speech-language therapy community.
Related Work
Recent research demonstrates the impressive performance of LLMs in zero-shot prompting across various tasks such as natural language inference, question answering, summarization, named-entity recognition, and sentiment analysis (Agrawal et al. 2022; Kojima et al. 2022; Qin et al. 2023; Zhong et al. 2023b). Considering LLMs’ limitations (Abid, Farooqi, and Zou 2021; Bang et al. 2023), some work explores the possibility of LLMs and humans working collaboratively to annotate the same dataset. However, mining structured knowledge from long-range unstructured text is much more challenging for human annotators due to the vast volume of content, the inherent ambiguity of natural language, and the individual bias of human interpretation (Ye et al. 2022). These challenges are further compounded by the need for specialized expertise, labor-intensive manual annotation, and difficulties in maintaining consistency and scalability.
Traditional methods for knowledge graph construction (KGC) rely on NLP techniques, using pipelines for named entity recognition and relation extraction (Martins, Marinho, and Martins 2019; Wei et al. 2019; Zhong et al. 2023a; Laurenzi, Mathys, and Martin 2024). Recent advancements leverage LLMs to generate relational triplets in zero/few-shot settings. For example, ChatIE (Wei et al. 2023) uses multi-turn zero-shot question-answering for triplets extraction, achieving promising results. Some studies (Zhang and Soh 2024; Carta et al. 2023) further automate KGC by decomposing it into several phases and allowing LLMs to infer knowledge graph schemas without predefined ontologies. However, these methods are typically limited to processing short texts and have not been tested in practical scenarios. For long-form, domain-specific texts, expert-crafted ontologies can guide LLMs extract expected knowledge (Cauter and Yakovets 2024). Our approach introduces a novel method for constructing domain-specific knowledge graphs automatically using LLMs, which we detail in the following sections.
Preliminaries
@@
Knowledge graph (KG) is a semantic network structured as an ontology, comprising concepts and their relationships with clear, interpretable structures at scale (Ehrlinger and Wöß 2016; Peng et al. 2023). For knowledge discovery from literature, LLMs can enhance this process by leveraging their capabilities to understand long-range text. This allows for transforming unstructured data into structured formats, and finally, populating the ontology to create KG.
@@
Focusing on knowledge discovery from literature, we aim to gather domain-specific knowledge for researchers and practitioners across various fields. In our approach, the ontology is predefined by domain experts, which can be represented by a directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{F})$. Here $\mathcal{E}$, $\mathcal{R}$, and $\mathcal{F}$ are sets of concepts, relationships, and semantic triples respectively. $\mathcal{F}$ is a collection of triples $(h,r,t)$ with a head concept $h\in\mathcal{E}$, a tail concept $t\in\mathcal{E}$, and a relation $r\in\mathcal{R}$ (Gruninger 1995). Named entity recognition (NER) and relation extraction (RE) are foundational elements for constructing knowledge graphs. To effectively instruct LLMs for knowledge discovery anchored to $\mathcal{G}$, the design of prompts and the sequencing of queries for both NER and RE tasks are essential for automating knowledge graph completion. We frame the problem of automated knowledge discovery via LLMs as one of prompt design and scheduling, described by the following equation:
$f(\mathcal{G}(\mathcal{E},\mathcal{R},\mathcal{F}))=\left\{(Prompt_{i},Order_{i})|i\in[1,N]\right\}$ (1)
where $f$ is a function that translates KG ontology into a set of prompt and action orders for the LLMs. A simple case of $f$ is directly injecting the whole ontology into a single prompt to instruct LLMs to generate KG annotation at once such as the annotation methods used in (Mihindukulasooriya et al. 2023; Kommineni, König-Ries, and Samuel 2024; Mihindukulasooriya et al. 2023).
Methodology
In this section, we first introduce the POP algorithm to convert a predefined knowledge ontology into a set of prompt templates and action orders, then propose an interactive annotation framework based on two LLM agents to enable more convincing and accurate annotation generations.
Ontology Prompting Algorithm
Figure 1: Illustration of prompt design and scheduling based on outdegree-prioritized BFS over a simple healthcare ontology for automating knowledge discovery.
@@
Algorithm 1 Progressive Ontology Prompting
1:function OntologyTraversal($\mathcal{G}$, $k$)
2:     Initialize: $queue,Order\leftarrow[]$, $visited,Khops\leftarrow\{\}$
3:     Identify the source node $s$
4:     $enqueue(queue,(s,0))$, $visited[s]\leftarrow\text{true}$, $Khops[s]\leftarrow[]$
5:     while not $queue$.empty() do
6:         $(current,hop)\leftarrow dequeue(queue)$
7:         $Order$.append$(current)$
8:         Gather unvisited neighbors of $current$ into $neighbors$
9:         Sort $neighbors$ by out-to-in ratio in descending order
10:         for $neighbor$ in $neighbors$ do
11:              if not $visited[neighbor]$ then
12:                  $enqueue(queue,(neighbor,hop+1))$
13:                  $visited[neighbor]\leftarrow\text{true}$
14:                  $Khops[neighbor]\leftarrow[]$
15:              end if
16:              if $hop+1\leq k$ then
17:                  $Khops[neighbor].append(current)$
18:              end if
19:         end for
20:     end while
21:     return $Order$, $Khops$
22:end function
23:function Main($\mathcal{G}$, $k$, $T$)
24:     Initialize $annotations\leftarrow\{\}$
25:     $Order,Khops\leftarrow\textsc{OntologyTraversal}(\mathcal{G},k)$
26:     while not $Order$.empty() do
27:         $current\leftarrow Order$.dequeue(), $context\leftarrow Khops[current]$
28:         if not $context$.empty() then
29:              $discoveries\leftarrow\{annotations[t]\text{ for }t\text{ in }context\}$
30:              $prompt\leftarrow T(current,G_{sub}(context),discoveries)$
31:         else
32:              $prompt\leftarrow T(current)$
33:         end if
34:         $annotations[current]\leftarrow\textsc{LLM}(prompt)$
35:     end while
36:end function
@@
In our work, we develop a progressive ontology prompting (POP) algorithm that employs a prioritized BFS on the ontology graph $\mathcal{G}(\mathcal{E},\mathcal{R},\mathcal{F})$ to generate a set of prompt templates and action orders for LLMs. As illustrated in Figure 1, in our method, the prompt design and scheduling follow a progressive manner. The annotation process begins by annotating the concept at a source node (i.e., nodes that only have outgoing edges) and then proceeds to its neighbors following the traversal order of a prioritized BFS. To allow BFS to quickly access a large portion of the graph, we modify BFS by sorting the neighbor nodes according to the out-to-in ratio $R(v)$, which is defined by:
$R(v)=\frac{|\left\{(h,r,t)\in\mathcal{F}\right\}|h=v|}{|\left\{(h,r,t)\in\mathcal{F}\right\}|t=v|}$ (2)
@@
Therefore, the algorithm will choose the neighbor node with higher $R$ to visit at the next step. For example, in Figure 1, visiting the ‘Patient’ node before the ‘Disorder’ node can provide more information for annotating the ‘Disorder’ concept. For any concept node $v$, we use the visited nodes within its k-hop neighborhood as its context. The $Prompt_{v}$ for identifying concept $v$ is crafted based on its own concept type and knowledge discoveries within its context. The action order of this prompt $Order_{v}$ is determined by the order in which node $v$ is visited during the prioritized BFS traversal.
@@
The pseudocode of the POP algorithm is shown in Algorithm 1. The algorithm first follows prioritized BFS traversal to capture the context and visit order of a specific concept node, then composes annotation prompts based on the local ontology structure and discoveries within its context as follows:
$Prompt(v)$$\leftarrow\left\{Prefix\left(N_{k-1}(u)\right)\oplus\right.$
$\left.Question\left((v,e,u)\mid(v,e,u)\in F\right)\mid u\in N_{1}(v)\right\}$ (3)
where $\oplus$ is the concatenation. We introduce two parameters in POP: context size $k$ and prompt template $T$. As the prompt examples shown in Figure 1, to annotate a specific concept, in prompt template $T$, we task LLMs to describe the knowledge discoveries within its context as a prefix to the annotation question associated with its concept type. Example prompts and corresponding outputs can be referred to in Appendix 3.
LLM-Duo Annotation Framework
Figure 2: Illustrative annotation process with LLM-Duo framework.
@@
Figure 3: Two annotation examples of speech-language intervention discovery using LLM-Duo framework. These examples demonstrate the adversarial annotation process with two LLMs when separately emphasizing on rationality and completeness.
To guarantee the integrity and reliability of knowledge discoveries via LLMs, we propose LLM-Duo. The architecture of LLM-Duo is illustrated in Figure 2. It comprises two primary LLM components: the explorer and the evaluator. Specifically, the explorer is a chatbot developed with RAG and works for the intervention discovery and annotation by zero-shot question answering (QA). RAG is utilized to ground its generations within the given literature context. Even though RAG can reduce the hallucination of LLMs by providing references to guarantee the integrity and reliability of explorers’ generations, the explorer might still make mistakes. To ensure maximized accuracy and reliability of annotations, we introduce another LLM, called the evaluator, tasked with inspecting the explorer’s responses to enhance the annotation quality.
In our annotation framework, the interaction between explorer and evaluator is collaborative and adversarial. Given a research paper, LLM-Duo will be tasked with prompts following a specific action order obtained from the POP algorithm. Depending on the exact requirements of annotations for different knowledge domains, the inspection mechanism of the evaluator can be customized. In our test case of speech-language intervention discovery, we emphasize the factual correctness and completeness of the annotations for different concepts, which are also typical requirements for annotation quality in most scenarios. LLMs cannot directly assess the correctness of the answer since they do not have prior knowledge of the correct response. Inspired by the concept of LLMs self-correcting through reasoning, we convert the correctness inspection to the answer’s rationality assessment.
In an interaction cycle of LLM-Duo for annotating concepts emphasizing rationality, the answer along with its explanatory reason is presented to the evaluator. The evaluator scrutinizes the answer’s rationality and gives feedback to the explorer. Depending on the feedback, the explorer may either adjust its answer accordingly or, if disagreeing, present stronger evidence to support its original answer to challenge the evaluator’s assessments. For annotating concepts emphasizing completeness, in each round, the evaluator will extract the aspects covered in the explorer’s answer and integrate them with aspects of the last round to form a new aspect collection and greedily prompt the explorer to delve deeper and explore more beyond the new aspect collection. The iterative refinement process will continue until the annotation achieves consistency. The key idea behind our LLM-Duo framework is using the feedback loop to assist the explorer refine its annotations. As an example shown in Figure 3, by facilitating interactive loops between two LLM agents, LLM-Duo enables more convincing and accurate annotations.
Experiments
Implementation

### v2

# Yuting Hu1    Dancheng Liu1    Qingyun Wang2    Charles Yu2    Chenhui Xu1    Qingxiao Zheng1    Heng Ji2&Jinjun Xiong1

@@ Abstract
Identifying effective interventions from the scientific literature is challenging due to the high volume of publications, specialized terminology, and inconsistent reporting formats, making manual curation laborious and prone to oversight. To address this challenge, this paper proposes a novel framework leveraging large language models (LLMs), which integrates a progressive ontology prompting (POP) algorithm with a dual-agent system, named LLM-Duo. On the one hand, the POP algorithm conducts a prioritized breadth-first search (BFS) across a predefined ontology, generating structured prompt templates and action sequences to guide the automatic annotation process. On the other hand, the LLM-Duo system features two specialized LLM agents, an explorer and an evaluator, working collaboratively and adversarially to continuously refine annotation quality. We showcase the real-world applicability of our framework through a case study focused on speech-language intervention discovery. Experimental results show that our approach surpasses advanced baselines, achieving more accurate and comprehensive annotations through a fully automated process. Our approach successfully identified 2,421 interventions from a corpus of 64,177 research articles in the speech-language pathology domain, culminating in the creation of a publicly accessible intervention knowledge base  [^1] Project website: https://slp.xlabub.com/ with great potential to benefit the speech-language pathology community.
1 Introduction
@@
Evidence-based interventions refer to practices and treatments grounded in systematic research and proven effective through controlled studies Rutten et al. (2021)Melnyk and Fineout-Overholt (2022). It emphasizes the use of evidence from well-designed and well-conducted research as the foundation for healthcare decision-making Sackett (1997). Intervention discovery from scientific literature enables researchers to keep abreast of the latest advancements and facilitate valuable insights that can significantly enhance the healthcare quality Usai et al. (2018)Wang et al. (2023a). However, due to the labor-intensive nature of human review, only a small fraction of intervention knowledge is systematically collected and curated. In healthcare, one of the biggest challenges for healthcare providers is the efficient identification of relevant intervention evidence from an overwhelming body of research, highlighting the urgent need for automated knowledge extraction tools to streamline the process and enhance the accessibility of this valuable information.
In recent years, large language models (LLMs) have been employed to categorize research papers, extract key findings, summarize complex studies, and create conversational assistants for question-answering and note generation, showing their impressive ability in understanding and extracting valuable insights from text Achiam et al. (2023)Li et al. (2024a). Many studies have utilized LLMs to streamline various subtasks involved in knowledge graph construction, such as named entity recognition (NER), relation extraction (RE), event extraction (EE), and entity linking (EL) Wang et al. (2023b)Zhu et al. (2024). Some research has also explored the collaboration between LLMs and human annotators to improve annotation quality Kim et al. (2024a)Wang et al. (2024)Tang et al. (2024). However, extracting intervention knowledge from long-range, domain-specific literature remains a significant challenge. On the one hand, developing human-annotated datasets for training deep learning models in NER and RE tasks requires specialized domain expertise to accurately interpret the literature. On the other hand, mining knowledge from long-range documents is a great challenge due to the vast volume of content, the inherent ambiguity of natural language, and the individual bias of human interpretation Ye et al. (2022). Particularly in healthcare contexts, these challenges are further compounded by the need for specialized therapeutic expertise, labor-intensive manual annotation, and difficulties in maintaining consistency and scalability Zhao et al. (2021). In this context, LLMs offer a promising alternative through in-context learning, enabling scalable information extraction without the need for extensive human-labeled data. Despite these advancements, fully automated knowledge graph construction remains a challenge, particularly when dealing with long-range documents. Most current knowledge graph construction approaches focus on short texts, leaving significant potential for further development in handling more complex, lengthy content.
@@
In this paper, we address the challenge of automating intervention discovery via LLMs by formulating it as a prompt design and annotation scheduling problem with a predefined intervention ontology graph structure and designing a framework leveraging two LLM agents to iteratively enhance the annotation quality. Specifically, we introduce a progressive ontology prompting (POP) algorithm that employs an outdegree-prioritized breadth-first search (BFS) across the intervention ontology to create a series of prompt templates and action sequences to guide the annotation process conducted by LLMs. To enhance the annotation quality, we propose LLM-Duo, an interactive annotation framework by leveraging the power of LLMs while addressing the limitations of LLMs. Particularly, it integrates two LLM agents working both collaboratively and adversarially to refine annotation generation.
@@
To showcase the practical impact of our approach, we apply our method in a case study of speech-language intervention discovery. We conduct experiments to compare our intervention discovery framework with several advanced baselines including long context LLM (i.e., GPT-4-Turbo with 128k context window length), and RAG-based annotation chatbot with advanced prompting techniques including Chain-of-Thought (CoT) Wei et al. (2022) and Self-Refine Madaan et al. (2024). The experimental results demonstrate that our method not only delivers more accurate and comprehensive annotations over these strong baselines but also significantly accelerates the intervention discovery process. Furthermore, through our framework, we successfully curate a speech-language intervention knowledge base, providing a valuable resource for the speech-language pathology community. To our knowledge, this is the first intervention knowledge base in the speech-language pathology field.
Related Work. Traditional approaches to automated knowledge discovery typically rely on pipelines to handle various NLP tasks such as named entity recognition, relation extraction, coreference resolution, entity linking, and event detection Luan et al. (2018)Martins et al. (2019)Wei et al. (2019)Zhong et al. (2023)Laurenzi et al. (2024). Recent advancements leverage LLMs to generate relational triplets in zero/few-shot settings for knowledge graph construction, achieving promising results Wei et al. (2023)Sun et al. (2024)He et al. (2024). Some studies Zhang and Soh (2024)Carta et al. (2023)Vamsi et al. (2024)Zhu et al. (2024) have further streamlined knowledge graph construction by breaking it down into distinct phases, enabling LLMs to infer knowledge graph schemas without relying on predefined ontologies. However, these methods are often constrained to short texts or have only been validated on tasks like entity and relation extraction using human-annotated datasets, such as DuIE2.0 Li et al. (2019) and DocRED Yao et al. (2019), without being proven effective in real-world applications. Moreover, domain-specific knowledge often exhibits complex patterns that cannot be captured solely through sentence-level syntactic structures. As a result, most existing approaches Du et al. (2020)Rossanez et al. (2020)Alam et al. (2023) are limited to handling abstracts and fail to extract and summarize knowledge across long-range contexts.
2 Preliminaries
@@
A knowledge graph (KG) is a semantic network structured as an ontology, consisting of concepts and their relationships in a clear, interpretable format at scale Peng et al. (2023). For intervention knowledge discovery from literature, LLMs can enhance this process by leveraging their capabilities to understand long-range text. This allows for transforming unstructured data into structured formats, and finally, populating the intervention ontology to create the intervention knowledge graph.
@@
In our methodology, the intervention KG ontology is crafted by domain experts, which can be represented by a directed acyclic graph (DAG) $\mathcal{G}=(\mathcal{E},\mathcal{R},\mathcal{F})$. Here $\mathcal{E}$, $\mathcal{R}$, and $\mathcal{F}$ are sets of concepts, relationships, and semantic triples respectively. $\mathcal{F}$ is a collection of triples $(h,r,t)$ with a head concept $h\in\mathcal{E}$, a tail concept $t\in\mathcal{E}$, and a relation $r\in\mathcal{R}$ Gruninger (1995). To effectively instruct LLMs to extract intervention knowledge anchored to $\mathcal{G}$ automatically, the design of annotation prompts and the query sequences plays a crucial role. We thereby frame the problem of automated intervention knowledge discovery via LLMs as one of prompt design and scheduling, described by the following equation:
$\small f(\mathcal{G}(\mathcal{E},\mathcal{R},\mathcal{F}))=\left\{(Prompt_{i},Order_{i})|i\in[1,N]\right\}$ (1)
where $f$ is a function that translates intervention KG ontology into a set of annotation prompts and query sequences for the LLMs. A common case of $f$ is directly prompting LLMs to generate triplets in a zero-shot/few-shot manner by including the whole KG schema within the prompt such as the annotation methods used in Mihindukulasooriya et al. (2023)Kommineni et al. (2024). However, those methods generate annotations in one shot and ignore the importance of contextual correlations between concepts within their surrounding neighborhood, resulting in incomplete annotations.
@@
Figure 1: Illustration of prompt design and scheduling based on the progressive ontology prompting algorithm.
3 Methodology
In this section, we first introduce the POP algorithm that converts an intervention KG ontology into a set of annotation prompt templates and query orders, then propose an interactive annotation framework based on two LLM agents to enable more convincing and accurate annotation generations.
3.1 Progressive Ontology Prompting
@@
We develop a progressive ontology prompting (POP) algorithm that employs a prioritized BFS on the intervention ontology graph $\mathcal{G}(\mathcal{E},\mathcal{R},\mathcal{F})$ to generate a set of annotation prompt templates and query sequences for LLMs. In our algorithm, the prompt formulation and scheduling follow a progressive manner. As illustrated in Figure 1, the annotation process begins at a source node (i.e., a concept node that only has outgoing edges) and continues by traversing its neighboring nodes in the order of a prioritized BFS. To allow for quick accessing a large portion of the graph, we enhance BFS by sorting neighboring nodes based on their out-to-in ratio $R(v)$, which is defined by:
$\small R(v)=\frac{|\left\{(h,r,t)\in\mathcal{F}\right\}|h=v|}{|\left\{(h,r,t)\in\mathcal{F}\right\}|t=v|}$ (2)
@@
Our algorithm selects the neighboring node with the maximum $R$ value to visit in the next step. For instance, in the example of Figure 1, visiting the ‘Patient’ node before the ‘Disorder’ node can provide more context for the ‘Disorder’ concept annotation. For each concept node $v$, we use the visited nodes within its $k$ hop neighborhood as its context. The $Prompt_{v}$ for annotating concept $v$ is crafted based on its context and the completed annotations within that context. The action order $Order_{v}$ for $Prompt_{v}$ is determined by the sequence in which node $v$ is visited during the prioritized BFS traversal.
@@
Our algorithm first follows prioritized BFS traversal to capture the local $k$ hop context and visit order of a specific concept node $v$, then composes the annotation prompts $Prompt_{v}$ based on its ontology substructure $N_{k}(v)$ and completed annotations within its context. This process can be expressed as follows:
$Prompt_{v}$ $\leftarrow T_{v}(Annotation(N_{k}(v)))$ (3)
$T_{v}$$\leftarrow\left\{Prefix\left(N_{k-1}(u)\right)\oplus\right.$
$\hskip 7.00002pt\left.Question\left((v,e,u)\mid(v,e,u)\in F\right)\mid u\in N_{1}(v)\right\}$ (4)
, where $\oplus$ is the concatenation. $Prompt_{v}$ represents a set of annotation prompts for node $v$, generated by applying completed annotations to the prompt template $T_{v}$. As illustrated in Figure 1, the prompt template $T_{v}$ consists of two parts: 1) Question, an annotation question derived from the relationship between node $v$ and one of its neighboring nodes $u$; and 2) Prefix, a description based on the $k-1$ hop path of neighbor node $u$. We leverage few-shot learning to task LLMs in generating prompt templates.
3.2 LLM-Duo Annotation Framework
Figure 2: Iterative annotation with two LLM agents under the LLM-Duo framework.
@@ 3.2 LLM-Duo Annotation Framework
Figure 3: Annotation examples of speech-language intervention discovery using the LLM-Duo framework.
To guarantee the integrity and reliability of LLM annotations, we propose LLM-Duo, a dual-agent annotation system. The central idea of a multi-agent system is to employ combinations of LLMs that can converse with each other to collaboratively accomplish tasks Wu et al. (2023). Drawing inspiration from the multi-agent debate idea in Kim et al. (2024b), we developed a framework where agents work both collaboratively and adversarially to enhance the quality of annotations.
The architecture of LLM-Duo is shown in Figure 2, featuring two LLM agents: the explorer and the evaluator. Specifically, the explorer is a chatbot performing annotation tasks using zero-shot question answering (QA). To break the context window limit of LLMs and ensure the generated annotations are faithful to the provided literature content, RAG is employed in explorer to reference relevant sources, minimizing LLM hallucinations. To further improve the accuracy and reliability of the annotations, the evaluator LLM is incorporated to review and validate the explorer’s responses, ensuring higher-quality results.
LLM-Duo will be tasked with annotation prompts following the sequential order generated by the POP algorithm. During each annotation cycle for a specific concept node, when focusing on concepts that emphasize rationality (e.g., disorder, intervention efficacy), the explorer provides an answer and an explanatory rationale to the evaluator. The evaluator then reviews the reasoning and offers feedback. Based on this feedback, the explorer either refines its answer or, if in disagreement, presents stronger evidence to defend its original answer and challenge the evaluator’s judgment. For concepts that emphasize completeness (e.g., intervention procedure, therapy activity), the evaluator extracts the aspects covered in each round of the explorer’s answer, combines them with aspects from previous rounds, and prompts the explorer to expand further beyond the newly integrated aspect collection. This iterative process continues until the annotations reach a consistent and comprehensive state. As the example shown in Figure 3, by facilitating interactive loops between two LLM agents, LLM-Duo enables more accurate and complete annotations.
4 Experiments
4.1 Implementation